OpenStudy (anonymous):

Solve the following equation for 0 degrees < x < 360 degrees 3 tan x + cot x = 3 sec x I will post my attempt at an answer.

OpenStudy (anonymous):

\[3tanx+\frac{1}{tanx}=3\sqrt{1+tan^{2}}\] multiply both sides by itself \[3tan^{2}x+6+\frac{1}{tan^{2}x}=9+9tan^{2}\]

OpenStudy (anonymous):

that should be \[9tan^{2}x+6+\frac{1}{tan^{2}x}=9+9tan^{2}\]

OpenStudy (anonymous):

\[6+\frac{1}{tan^{2}x}=9\] \[\frac{1}{tan^{2}x}=3\] \[tan^{2}x=\frac{1}{3}\] \[tanx=\frac{1}{+-\sqrt3}\]

OpenStudy (anonymous):

x = 30, 210, 150, 330 The answer my textbook gives is 30, 90, 150 Help?

OpenStudy (eyust707):

"multiply both sides by itself"

OpenStudy (eyust707):

thats where you went wrong I beleive

OpenStudy (eyust707):

You did not correctly square the left side

OpenStudy (anonymous):

Sorry but did you check the very next post.

OpenStudy (eyust707):

nope stopped when i hit the mistake =P

OpenStudy (eyust707):

I shall continue

OpenStudy (eyust707):

@amistre64 ?

OpenStudy (amistre64):

i tend to put everything into sins and coss

OpenStudy (anonymous):

Maybe I'll get the book's answer if I try that. It is still frustrating if I don't know what is wrong with what happened here though. I can definitely see a way to solve this with sins and coss now that you mention it.

OpenStudy (amistre64):

\[3\frac{s}{c}+\frac{c}{s}=3\frac{1}{c}\] \[3s+\frac{cc}{s}=3\] \[3s+\frac{1-ss}{s}=3\] \[\frac{3ss+(1-ss)}{s}=\frac{3s}{s}\] \[2s^2+1=3s\] \[2s^2-3s+1=0\] might be one option to try

OpenStudy (anonymous):

I just solved it using your suggestion and it matches. Thanks. Would be nice to know the problem with my initial attempt though.

OpenStudy (eyust707):

Possible domain issue??

OpenStudy (amistre64):

3t*3t not= 3t^2 is one error (possibly typo?)

OpenStudy (amistre64):

you corrected with 2nd post

OpenStudy (amistre64):

if i were to take a guess, i would say that squaring should be a last resort since it has a tendency to introduce extraneous solutions at times

OpenStudy (anonymous):

Well thanks for the help.

OpenStudy (amistre64):

good luck ;)